OpenStudy (anonymous):

Given v = (4,9) a. What is the angle between the vector and the x-axis? b. Write the vector in polar form

OpenStudy (amistre64):

whats its tangent inverse?

OpenStudy (amistre64):



OpenStudy (amistre64):

polar form: (length,angle) you know the angle, whats its length?

OpenStudy (anonymous):

\[\sqrt{4^{2}+9^{2}}\] = \[\sqrt{97}\] is what I have

OpenStudy (anonymous):

The length based on that triangle you drew, using the Pythagorean theorem would make the length equal to 97, right?

OpenStudy (anonymous):

sqt rt of 97 *

OpenStudy (anonymous):

so the length would be 9.848857802

OpenStudy (amistre64):

sqrt 97 is a good length, dunno how 'simplified' it is but its valid to me

OpenStudy (anonymous):

so I have the length, I need the angle now. How could I set that up?

OpenStudy (amistre64):

already posted it .... tan(angle) = 9/4 take the inverse ...

OpenStudy (anonymous):

66.03751103

OpenStudy (amistre64):

even if we do a dot product equivalence, we still have to do an inverse so might as well work it from the start lol

OpenStudy (amistre64):

yeah, about 66 degrees

OpenStudy (amistre64):

exactness is: (sqrt(97), arctan(9/4)) rnd them to whatever you like if need be

OpenStudy (anonymous):

(9.849, 66.038) yeah I need to do it in at least the thousandths place

OpenStudy (amistre64):

:) ok

OpenStudy (anonymous):

So now that I have the length and angle, how can i set that up in polar formation?

OpenStudy (anonymous):

wait, is that in polar form already? (9.849, 66.038)? @amistre64